# Two-interface and thin filament approximation in Hele–Shaw channel flow

### v2

@@ 5 Numerical simulation and comparison
In this section we describe numerical simulations of the thin filament model (3), and compare against simulations of the full two-interface problem (2.1).
@@ 5.1 Validation against solution of the two-interface problem
To validate solutions of the thin filament model (3) it is valuable to compare it to solutions of the full two-interface system (2.1), which is a more challenging numerical problem. We solve (2.1) using the numerical framework proposed by Morrow et al. (2021, 2023), which we briefly summarise here. The framework is based on the level set method, in which we represent each interface, $f_{\mathrm{L}}$ and $f_{\mathrm{U}}$, as the zero level set of the associated level set functions $\psi_{\mathrm{L}}$ and $\psi_{\mathrm{U}}$. Each level set function is chosen such that the viscous fluid will occupy the region where both $\psi_{\mathrm{L}}$ and $\psi_{\mathrm{U}}>0$; otherwise the region is filled with inviscid fluid. Both level set functions are updated according to
@@ 5.2 Numerical results for late times
Here we use the front-tracking scheme to solve the thin filament model for later times, in the regime where the thickness becomes too small for the level set method to accurately resolve. Here we use the initial condition
@@ 6 Discussion
In this paper we have developed a simplified but highly accurate second-order lubrication flow model (3) that describes two-interface Hele–Shaw flow very well in regions where the thickness of the fluid region becomes small. Due to the instability of one of the interfaces, this limit is one that is generally reached, even if initially the thickness is not very small. By examining the generic singular behaviour of the leading-order model, even in the presence of surface tension (Section 4), we have shown why a second order-model is necessary to represent the dynamics of the full problem.

### v3

@@ 5 Numerical simulation and comparison
In this section we describe numerical simulations of the second-order thin filament model (16), and compare against simulations of the full two-interface problem (1).
@@ 5.1 Validation against solution of the two-interface problem
To validate solutions of the thin filament model (16) it is valuable to compare it to solutions of the full two-interface system (1), which is a more challenging numerical problem. We solve (1) using the numerical framework proposed by Morrow et al. (2021, 2023), which we briefly summarise here. The framework is based on the level set method, in which we represent each interface, $f_{\mathrm{L}}$ and $f_{\mathrm{U}}$, as the zero level set of the associated level set functions $\psi_{\mathrm{L}}$ and $\psi_{\mathrm{U}}$. Each level set function is chosen such that the viscous fluid will occupy the region where both $\psi_{\mathrm{L}}$ and $\psi_{\mathrm{U}}>0$; otherwise the region is filled with inviscid fluid. Both level set functions are updated according to
@@ 5.2 Numerical results for late times
Here we use the front-tracking scheme to solve the thin filament model for later times, in the regime where the thickness becomes too small for the level set method to accurately resolve. We use the initial condition
@@
5.3 Initial condition near a quasi-travelling wave
In addition to the previous near-flat initial conditions, we demonstrate the instability of the quasi-travelling wave solutions considered in Section 4.3 by choosing an initial condition close to one such solution. For simplicity we use the ‘grim reaper’ exact solution for zero surface tension $K_{0}=-\cos\theta$, and a scale factor $B=1$, for which
$y=\log(\cos x),\qquad h=\frac{b}{\cos x}$
(the exact solution is a good approximation for small $\sigma$, and although we do not show the results here, we have observed the same behaviour from numerically constructed initial conditions by solving (32) for $\hat{\sigma}>0$). In terms of the quasi-travelling wave solution, the factor $b$ in the thickness $h$ is arbitrary as it corresponds to translation in time. As the travelling wave is semi-infinite, we must of course approximate it by a finite curve that respects the required boundary conditions at $x=\pm\pi$. We thus choose the closely related ‘hairclip’ curve, defined by
$y=\sinh^{-1}(\alpha\cos x),\qquad h=\frac{b}{\cos\theta}=b\sqrt{1+\left(\frac{\alpha\sin x}{\cosh y}\right)^{2}},\qquad\alpha\gg 1.$ (40)
The larger the parameter $\alpha$ is made, the more elongated the initial finger becomes. We note that the thickness will become large on the sides of the initial finger, representing the blow-up in thickness of the quasi-travelling wave in the limit $\theta\to\pi/2$ discussed in Section 4.3.
We depict the result of such an initial condition in Fig. 10, for the values $\sigma=0.05$, $\alpha=20$, $b=0.1$. The regions with large thickness on the sides of the finger correspond to the exponential growth in thickness required for the travelling wave solution as $y\to-\infty$ (see Section 4.3); the filament in this region does not evolve to a great extent over the simulation. At the front, while the finger does initially propagate forwards, the nose bulges outward and becomes more circular in shape, leading to the same late-time behaviour as seen for the more general initial conditions.
Figure 10: The evolution of a filament that starts near the ($\sigma=0$ limit) of the quasi-travelling wave described in Section 4.3 (the precise initial condition is (40)). (a) the quasi-travelling wave appears unstable, with the front evolving toward the expanding circle seen for general initial conditions. (b) The curvature at the nose of the filament ($x=0$) starts at $-1$ but decreases over time as the filament expands. The dots marked in (b) correspond to the profiles plotted in (a).
@@ 6 Discussion
In this paper we have developed a simplified but highly accurate second-order lubrication flow model (16) that describes two-interface Hele–Shaw flow very well in regions where the thickness of the fluid region becomes small. Due to the instability of one of the interfaces, this limit is one that is generally reached, even if initially the thickness is not very small. By examining the generic singular behaviour of the leading-order model, even in the presence of surface tension (Section 4), we have shown why a second order-model is necessary to represent the dynamics of the full problem. In particular, the fourth order spatial derivative term arising from the difference in curvature between the upper and lower interfaces is necessary to regularise the system. Although we have included all terms formally of order $\epsilon^{2}$ in our model and simulations, we have also observed that the leading order model along with the addition of only this regularising term (proportional to $[hh_{sss}]_{s}$) will behave in a qualitatively similar manner (with small quantitative differences). We have not shown these results here for the sake of brevity.